OpenStudy (anonymous):

find the equation of the curve passing through the point (0,2) and having the gradient e^x at p(x,y).

OpenStudy (anonymous):

\[y=e^{x }+1\]

OpenStudy (anonymous):

Let me wright you how...

OpenStudy (anonymous):

it is right but how can you tell me in detail

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=e^{x}\]\[dy = e^{x}.dx\]\[\int\limits dy = \int\limits e^{x}.dx\]\[y = e^{x}+c\]But it passes through (0, 2) ; thus c = 1 finally\[y=e^{x} + 1\]

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks